OpenStudy (anonymous):

@johnweldon1993 are you good at these??? Isabel bought a desktop computer and a laptop computer. Before finance charges, the laptop cost 250 more than the desktop. She paid for the computers using two different financing plans. For the desktop the interest rate was 8.5% per year, and for the laptop it was 5% per year. The total finance charges for one year were 296 . How much did each computer cost before finance charges?

OpenStudy (johnweldon1993):

I'm trying to do it out hang on :)

OpenStudy (johnweldon1993):

Ha yes got it...okay ready? you'll be doing the work :)

OpenStudy (johnweldon1993):

*p.s It's just another system of equations...as you'll see*

OpenStudy (johnweldon1993):

So step 1...have you gotten any equations written down from this information?

OpenStudy (anonymous):

i figured it out but i cant figure out this one A theater group made appearances in two cities. The hotel charge before tax in the second city was 1500 higher than in the first. The tax in the first city was 5% , and the tax in the second city was 7.5% . The total hotel tax paid for the two cities was $737.5 . How much was the hotel charge in each city before tax?

OpenStudy (johnweldon1993):

haha aww after all that work :P and this would be the same exact process as that computer one...

OpenStudy (anonymous):

okayy so y=x+1500 and 0.05x+0.75(x+1500)=737.5 ????

OpenStudy (johnweldon1993):

exactly

OpenStudy (johnweldon1993):

well no...not EXACTLY!!! lol....what is 7.5% in decimal form?

OpenStudy (anonymous):

.75????

OpenStudy (johnweldon1993):

you had it right with 5%.....you move the DECIMAL place over 2 to the left right? Just like in the computer problem...with the 8.5%...would be .085

OpenStudy (anonymous):

ohhh okay so its .875x=625??

OpenStudy (johnweldon1993):

woh....where did THOSE numbers come from? you have *with corrected percent "okayy so y=x+1500 and 0.05x+0.075(x+1500)=737.5 ????"

OpenStudy (anonymous):

i caught that lol so its 0.125x=625??

OpenStudy (anonymous):

so x = 5000 and y = 6000

OpenStudy (johnweldon1993):

Looks correct to me :) as long as you checked them

OpenStudy (johnweldon1993):

WAIT! no! lol...check them!!!

OpenStudy (anonymous):

i did!!!! lol is this the same way or different set up?? Henry bought a desktop computer and a laptop computer. Before finance charges, the laptop cost $200 less than the desktop. He paid for the computers using two different financing plans. For the desktop the interest rate was 5% per year, and for the laptop it was 7% per year. The total finance charges for one year were $250 . How much did each computer cost before finance charges?

OpenStudy (johnweldon1993):

"the second city was 1500 higher than in the first" YOU CHECKED THEM AND IT WORKED!?

OpenStudy (anonymous):

yeah and it was right lol

OpenStudy (johnweldon1993):

you used 5000 for x and 6000 for y???

OpenStudy (anonymous):

yeah lol

OpenStudy (johnweldon1993):

lmao please check it again! xD I'm so confused! haha

OpenStudy (anonymous):

6500 because 5000+1500=6500

OpenStudy (johnweldon1993):

HA! yes there we go! lol....you said 6000 before! haha

OpenStudy (anonymous):

oh stupid keyboard lol can u help with the new one??? lol and 1 more of this kind after

OpenStudy (johnweldon1993):

and this previous one you posted SAME THING...just remember "the laptop cost $200 less than the desktop"

OpenStudy (anonymous):

so y =200-x??

OpenStudy (johnweldon1993):

correct!

OpenStudy (anonymous):

and 0.05x-0.07y=250

OpenStudy (johnweldon1993):

so y =200-x?? should be y = x - 200...but yes...and yes on the percents

OpenStudy (anonymous):

so then is 0.05x+0.07x-14=250??

OpenStudy (johnweldon1993):

and 0.05x "+" 0.07y=250

OpenStudy (anonymous):

so then x=2200 and y = 2200-200 so y =2000

OpenStudy (johnweldon1993):

and check themmmmm lol *going to be my catch phrase :D

OpenStudy (anonymous):

lol i did and they were right :D you make this easy

OpenStudy (johnweldon1993):

We'll You're making it Easy lol...you're doing all the work :D

OpenStudy (anonymous):

you are helping lol okay last one of this kind... Last year, Chris had 10,000 to invest. He invested some of it in an account that paid simple 7% interest per year, and he invested the rest in an account that paid simple 5% interest per year. After one year, he received a total of $560 in interest. How much did he invest in each account?

OpenStudy (anonymous):

y=10,000-x and 0.07x+0.05y=560???

OpenStudy (johnweldon1993):

I'm going to let you do this one...you've surpassed my teaching :D

OpenStudy (anonymous):

x=3000 and y =7000??

OpenStudy (johnweldon1993):

Bravo! :)

OpenStudy (anonymous):

i checked and they were right!!!!!

OpenStudy (anonymous):

r u good at these Two buses leave towns 616mi apart at the same time and travel toward each other. One bus travels 8mi/h slower than the other. If they meet in 4 hours, what is the rate of each bus?

OpenStudy (johnweldon1993):

Pretty much still system of equations They meet in 4 hours...so 4x would represent the distance covered by the faster car here. So bus 1....travels 8mi/h slower.... x - 8 is the rate of bus 1 4(x-8) would be the rate of the slower car They are 616 mi apart so 4(x - 8) + 4x = 616 And solve :)

OpenStudy (johnweldon1993):

*sorry didn't really know how to explain it...so I did some of it lol

OpenStudy (anonymous):

x=81 so where do i plug it in at

OpenStudy (anonymous):

so rate one is 73 and the other is 292??

OpenStudy (johnweldon1993):

This is the rate of the faster car here....so the slower car would be that - 8

OpenStudy (anonymous):

so if the faster car is 81-8 its 73 so the other one is 4(73) so 292

OpenStudy (johnweldon1993):

so faster car rate is 81 and slower car rate is 81 - 8 = 73 and we can check by 4(x) + 4(y) = 616 4(81) + 4(73) = 616

OpenStudy (johnweldon1993):

*what you're doing...is solving how far the cars are going...they are asking for the rates only

OpenStudy (johnweldon1993):

buses*

OpenStudy (anonymous):

ohh i see

OpenStudy (johnweldon1993):

I feel like that one confused you??

OpenStudy (anonymous):

it did but it was right lol

OpenStudy (johnweldon1993):

Because I can probably make it easier to follow if I added another variable if you want?

OpenStudy (anonymous):

nope i got it lol

OpenStudy (johnweldon1993):

Okay, but I can :) lol

OpenStudy (anonymous):

lol if i need help with my final topic ill let u know lo

OpenStudy (johnweldon1993):

lol okay!